# Intrinsic Saliency Guided Trunk-Collateral Network for Unsupervised Video Object Segmentation

### v1

@@ IV-A3 Datasets
We provide an introduction of the datasets utilized in our work. ISTC-Net is pre-trained on the YouTube-VOS dataset and fine-tuned for specific tasks. For the UVOS task, the model is evaluated on the test sets of DAVIS-16, FBMS, and YouTube-Objects. For the VSOD task, we evaluate the performance of ISTC-Net on the test sets of DAVIS-16, FBMS, ViSal, and DAVSOD.
@@ IV-B1 Unsupervised Video Object Segmentation Results
To assess the effectiveness of ISTC-Net on the UVOS task, we fine-tuned and tested ISTC-Net on the DAVIS-16 and FBMS dataset. For the YouTube-Objects dataset, we directly leveraged the pre-trained model for testing. Results of ISTC-Net on UVOS metrics are presented in Table I, with comparison to past UVOS works. In particular, for studies reporting multiple results, we only present those results derived using standard evaluation protocols and widely adopted backbones.
@@
These prior approaches ignored the fundamental motion-appearance relationship, which resulted in the marginal improvements. In contrast, the trunk-collateral architecture of ISTC-Net accounts for motion-appearance distinctions and commonalities simultaneously, which facilitates feature encoding without relying on complex fusion strategies. Additionally, some methods, such as [46], incorporated additional inputs (e.g., depth maps or external saliency cues) to aid in segmentation. While effective to some extent, this fashion severely impedes the model’s efficiency and incurs greater storage costs. ISTC-Net instead leverages the model’s intrinsic saliency, which promotes model performance while streamlining the overall process.
@@
Ours(SRTC) 85.9 86.4 74.3 86.5 82.3 70.8 83.1 67.1 47.9 74.1 76.0
A series of outstanding experimental results demonstrate the superiority and effectiveness of our model. ISTC-Net achieves state-of-the-art performance across all three standard UVOS benchmarks. On the DAVIS-16 validation set, ISTC-Net attained 89.2% on the $\mathcal{J}\&\mathcal{F}$ metric, outperforming all previous models by a margin of at least 1%. On the FBMS test set, our model recorded a $\mathcal{J}$ score of 86.4% , marking a 3.0% improvement over the best recent results. Moreover, ISTC-Net outperformed all existing UVOS methods on the YouTube-Objects dataset, achieving a 1.3% $\mathcal{J}$ value improvement over the recent top-performing approach. As detailed in Table II, our model secured either the best or second-best scores in six out of all ten categories in the YouTube-Objects dataset, underscoring the exceptional efficacy of ISTC-Net.
@@ IV-B2 Video Salient Object Detection Results
Given the shared objective of detecting visually salient objects in a video sequence between UVOS and VSOD tasks, we also evaluate the performance of ISTC-Net on four VSOD benchmarks: DAVIS-16, FBMS, ViSal, and DAVSOD. To demonstrate the effectiveness of our model on these benchmarks, we employ four evaluation metrics and compare ISTC-Net’s results against those of previous UVOS methods, which either reported relevant results or provided publicly available code. The comparative results are presented in Table III.
@@
Our ISTC-Net consistently outperforms existing UVOS algorithms, achieving state-of-the-art performance across all four VSOD benchmarks. Of the total 16 evaluation metrics used across these datasets, ISTC-Net secured the best on 15 metrics and ranked second on the remaining one. Notably, ISTC-Net achieved the highest scores for the $F_{m}$, $E_{m}$, and $S_{m}$ metrics across all benchmarks with substantial improvements over prior methods, while attaining either the best or second-best performance on the $MAE$ metric. These outstanding quantitative results underscore the robustness and effectiveness of ISTC-Net in identifying salient objects in video sequences, further demonstrating its superior capability in addressing VSOD and UVOS tasks.
@@
Figure 6: Qualitative demonstration of ISTC-Net’s performance on diverse complex scenes from the DAVIS-16, FBMS, and YouTube-Objects datasets. The selected scenes include various challenging situations for the UVOS task, such as rapid object motion, co-moving background elements, multiple targets, intricate details, motion blur, and so on. Case 1-3: dance-twirl, drift-chicane, soapbox from DAVIS-16; Case 4-6: dogs02, cars4, rabbits02 from FBMS; Case 7-9: boat0012, car0009, horse0011 from YouTube-Obects.
@@ IV-C1 Unsupervised Video Object Segmentation
We qualitatively demonstrate the visual effectiveness of ISTC-Net on the UVOS task in Fig. 6. Nine representative scenes were selected from three widely used UVOS benchmarks: DAVIS-16, FBMS, and YouTube-Objects. These videos cover a diverse range of challenging UVOS scenarios, including rapid object motion, co-moving background elements, multiple targets, intricate details, motion blur, etc.
@@
Some videos involve objects undergoing substantial visual changes in size due to shifting positions. For example, in the second, fifth, and eighth cases, the target objects transition from appearing small to large as their distance from the camera varies. Despite these volume transitions, ISTC-Net maintains highly accurate segmentation, demonstrating its adaptability to such challenges.
Furthermore, in the second and seventh cases, the presence of water ripples and smoke introduced by the movement of the primary objects is a common source of interference in UVOS tasks. These dynamic elements, which often move in sync with the background, can compromise the performance of optical flow in distinguishing foreground targets. Nevertheless, ISTC-Net effectively isolates the foreground objects from the background, preserving precise contours.
Additionally, ISTC-Net successfully handles a variety of other complex scenarios, such as multiple targets, motion blur, and the emergence of new objects, as shown in selected cases. These results highlight the robustness and exceptional segmentation capabilities of ISTC-Net, even in highly challenging conditions.
@@ IV-C2 Video Salient Object Detection
We further illustrate the visual results of ISTC-Net in recognizing video salient objects on three challenging videos selected from the VSOD and FBMS datasets. The saliency detection maps are presented in Fig. 7, alongside predictions from prior UVOS algorithms for qualitative comparison.
The first video represents a low-light scenario, where previous methods often struggled to consistently distinguish the entire target across multiple frames. ISTC-Net accurately identifies salient objects even in dimly lit frames, and with significantly higher precision. The second video involves complex human motion and intricate details, posing significant challenges for the VSOD task. Prior approaches frequently produced incomplete detections or identified targets with low confidence. However, ISTC-Net precisely detects both individuals with high confidence, accurately capturing their limbs and other fine-grained details. In the third video, the target object undergoes drastic changes in shape and size. Previous methods either failed to consistently track the target or detected irrelevant objects. In comparison, ISTC-Net maintains high detection accuracy, reliably segmenting the target without interference.
The visual results presented in Fig. 7 highlight the remarkable performance of ISTC-Net across various complex scenarios. These findings emphasize the model’s superior capability to handle diverse challenges in VSOD with robustness and accuracy.
@@ IV-D Ablation Study
We performed a series of ablation studies to provide insights into the inner workings of ISTC-Net and validate its effectiveness. The conducted experiments focus on four aspects: the effectiveness of the trunk-collateral structure and ISRM; the contribution of appearance and motion features; the influence of the LoRA rank value; and the placement of the collateral branch within the Transformer block. All experiments are carried out on the DAVIS-16 dataset using the mit-b1 [56] backbone.
@@ IV-D1 Effectiveness of Different Modules
These results confirm the effectiveness of both the trunk-collateral structure and ISRM, underscoring their critical role in the success of ISTC-Net and validating the overall pattern of our proposed framework.
@@ V Conclusion
We present ISTC-Net, an innovative end-to-end UVOS framework, which better adapts to the motion-appearance intrinsic relationship while utilizing the model’s inherent saliency to enhance performance. ISTC-Net advances a novel trunk-collateral paradigm that regards both the motion-appearance correspondence and uniqueness. Additionally, we put forward the intrinsic saliency guided refinement module, which leverages the model’s inherent saliency to optimize high-level representations and guide the motion-appearance integration. Comprehensive experiments demonstrate the superiority of ISTC-Net, which achieved the state-of-the-art results on all commonly used UVOS and VSOD benchmarks. We hope that our model will provide some insights for better utilization of motion and appearance features and inspire future works.

### v2

@@ IV-A3 Datasets
We provide an introduction of the datasets utilized in our work. SMTC-Net is pre-trained on the YouTube-VOS dataset and fine-tuned for specific tasks. For the UVOS task, the model is evaluated on the test sets of DAVIS-16, FBMS, and YouTube-Objects. For the VSOD task, we evaluate the performance of SMTC-Net on the test sets of DAVIS-16, FBMS, ViSal, and DAVSOD.
@@ IV-B1 Unsupervised Video Object Segmentation Results
To assess the effectiveness of SMTC-Net on the UVOS task, we fine-tuned and tested SMTC-Net on the DAVIS-16 and FBMS dataset. For the YouTube-Objects dataset, we directly leveraged the pre-trained model for testing. Results of SMTC-Net on UVOS metrics are presented in Table I, with comparison to past UVOS works. In particular, for studies reporting multiple results, we only present those results derived using standard evaluation protocols and widely adopted backbones.
@@
These prior approaches ignored the fundamental motion-appearance relationship, which resulted in the marginal improvements. In contrast, the trunk-collateral architecture of SMTC-Net accounts for motion-appearance distinctions and commonalities simultaneously, which facilitates feature encoding without relying on complex fusion strategies. Additionally, some methods, such as [46], incorporated additional inputs (e.g., depth maps or external saliency cues) to aid in segmentation. While effective to some extent, this fashion severely impedes the model’s efficiency and incurs greater storage costs. SMTC-Net instead leverages the model’s intrinsic saliency, which promotes model performance while streamlining the overall process.
@@
Ours(SMTC-Net) 85.9 86.4 74.3 86.5 82.3 70.8 83.1 67.1 47.9 74.1 76.0
A series of outstanding experimental results demonstrate the superiority and effectiveness of our model. SMTC-Net achieves state-of-the-art performance across all three standard UVOS benchmarks. On the DAVIS-16 validation set, SMTC-Net attained 89.2% on the $\mathcal{J}\&\mathcal{F}$ metric, outperforming all previous models by a margin of at least 1%. On the FBMS test set, our model recorded a $\mathcal{J}$ score of 86.4% , marking a 3.0% improvement over the best recent results. Moreover, SMTC-Net outperformed all existing UVOS methods on the YouTube-Objects dataset, achieving a 1.3% $\mathcal{J}$ value improvement over the recent top-performing approach. As detailed in Table II, our model secured either the best or second-best scores in six out of all ten categories in the YouTube-Objects dataset, underscoring the exceptional efficacy of SMTC-Net.
@@ IV-B2 Video Salient Object Detection Results
Given the shared objective of detecting visually salient objects in a video sequence between UVOS and VSOD tasks, we also evaluate the performance of SMTC-Net on four VSOD benchmarks: DAVIS-16, FBMS, ViSal, and DAVSOD. To demonstrate the effectiveness of our model on these benchmarks, we employ four evaluation metrics and compare SMTC-Net’s results against those of previous UVOS methods, which either reported relevant results or provided publicly available code. The comparative results are presented in Table III.
@@
Our SMTC-Net consistently outperforms existing UVOS algorithms, achieving state-of-the-art performance across all four VSOD benchmarks. Of the total 16 evaluation metrics used across these datasets, SMTC-Net secured the best on 14 metrics and ranked second on the remaining one. Notably, SMTC-Net achieved the highest scores for the $F_{m}$, $E_{m}$, and $S_{m}$ metrics across all benchmarks with substantial improvements over prior methods, while attaining either the best or second-best performance on the $MAE$ metric. These outstanding quantitative results underscore the robustness and effectiveness of SMTC-Net in identifying salient objects in video sequences, further demonstrating its superior capability in addressing VSOD and UVOS tasks.
@@
Figure 6: Qualitative demonstration of SMTC-Net’s performance on diverse complex scenes from the DAVIS-16, FBMS, and YouTube-Objects datasets. The selected scenes include various challenging situations for the UVOS task, such as rapid object motion, co-moving background elements, multiple targets, intricate details, motion blur, and so on. Case 1-3: dance-twirl, drift-chicane, soapbox from DAVIS-16; Case 4-6: dogs02, cars4, rabbits02 from FBMS; Case 7-9: boat0012, car0009, horse0011 from YouTube-Obects.
@@ IV-C1 Unsupervised Video Object Segmentation
We qualitatively demonstrate the visual effectiveness of SMTC-Net on the UVOS task in Fig. 6. Nine representative scenes were selected from three widely used UVOS benchmarks: DAVIS-16, FBMS, and YouTube-Objects. These videos cover a diverse range of challenging UVOS scenarios, including rapid object motion, co-moving background elements, multiple targets, intricate details, motion blur, etc.
@@
Some videos involve objects undergoing substantial visual changes in size due to shifting positions. For example, in the second, fifth, and eighth cases, the target objects transition from appearing small to large as their distance from the camera varies. Despite these volume transitions, SMTC-Net maintains highly accurate segmentation, demonstrating its adaptability to such challenges.
Furthermore, in the second and seventh cases, the presence of water ripples and smoke introduced by the movement of the primary objects is a common source of interference in UVOS tasks. These dynamic elements, which often move in sync with the background, can compromise the performance of optical flow in distinguishing foreground targets. Nevertheless, SMTC-Net effectively isolates the foreground objects from the background, preserving precise contours.
Additionally, SMTC-Net successfully handles a variety of other complex scenarios, such as multiple targets, motion blur, and the emergence of new objects, as shown in selected cases. These results highlight the robustness and exceptional segmentation capabilities of SMTC-Net, even in highly challenging conditions.
@@ IV-C2 Video Salient Object Detection
We further illustrate the visual results of SMTC-Net in recognizing video salient objects on three challenging videos selected from the VSOD and FBMS datasets. The saliency detection maps are presented in Fig. 7, alongside predictions from prior UVOS algorithms for qualitative comparison.
The first video represents a low-light scenario, where previous methods often struggled to consistently distinguish the entire target across multiple frames. SMTC-Net accurately identifies salient objects even in dimly lit frames, and with significantly higher precision. The second video involves complex human motion and intricate details, posing significant challenges for the VSOD task. Prior approaches frequently produced incomplete detections or identified targets with low confidence. However, SMTC-Net precisely detects both individuals with high confidence, accurately capturing their limbs and other fine-grained details. In the third video, the target object undergoes drastic changes in shape and size. Previous methods either failed to consistently track the target or detected irrelevant objects. In comparison, SMTC-Net maintains high detection accuracy, reliably segmenting the target without interference.
The visual results presented in Fig. 7 highlight the remarkable performance of SMTC-Net across various complex scenarios. These findings emphasize the model’s superior capability to handle diverse challenges in VSOD with robustness and accuracy.
@@ IV-D Ablation Study
We performed a series of ablation studies to provide insights into the inner workings of SMTC-Net and validate its effectiveness. The conducted experiments focus on four aspects: the effectiveness of the trunk-collateral structure and ISRM; the contribution of appearance and motion features; the influence of the LoRA rank value; and the placement of the collateral branch within the Transformer block. All experiments are carried out on the DAVIS-16 dataset using the mit-b1 [56] backbone.
@@ IV-D1 Effectiveness of Different Modules
These results confirm the effectiveness of both the trunk-collateral structure and ISRM, underscoring their critical role in the success of SMTC-Net and validating the overall pattern of our proposed framework.
@@ V Conclusion
We present SMTC-Net, an innovative end-to-end UVOS framework, which better adapts to the motion-appearance intrinsic relationship while utilizing the model’s inherent saliency to enhance performance. SMTC-Net advances a novel trunk-collateral paradigm that regards both the motion-appearance correspondence and uniqueness. Additionally, we put forward the intrinsic saliency guided refinement module, which leverages the model’s inherent saliency to optimize high-level representations and guide the motion-appearance integration. Comprehensive experiments demonstrate the superiority of SMTC-Net, which achieved the state-of-the-art results on all commonly used UVOS and VSOD benchmarks. We hope that our model will provide some insights for better utilization of motion and appearance features and inspire future works.